Kip is trying to decide whether to take a one-day martial arts class at Rex Kwan Do. The fee for the class is $400. Kip will also have to pay his brother Napoleon $150 for that day in return for giving him a ride into town. Instead of taking the class Kip could spend all day on the computer chatting with his girl, LaFawnduh. This gives him a value of $300. On the other hand, he could spend all day working with his Uncle Rico and making sweet moolah. Uncle Rico has promised to pay him $350 on that day if he works with him. What is Kip’s marginal cost of joining Rex Kwan Do?
Answer #1

Here Kip's marginal cost is actually Kip's marginal opportunity cost.

Marginal opportunity cost = Sacrifice/ gains.

So here, gains = value received from chatting with ha girl + pay from his uncle.

Gains = 300 + 350 = 650

Sacrifice = value he will be spending to obtain the classes.

= Fees + pay to his brother

= 400 + 150 = 550

MOC = 550/650 = 0.846.

So, marginal cost of doing Rex Mean Do is 0.85 units.